OpenStudy (anonymous):

With the restriction x ≠ 0, Which of the following is the simplified form of 6x^4/4x^9 x 12x^2/3x^5

OpenStudy (paxpolaris):

\[{6x^4\over \cancel4x^9} \times {\cancel3\cancel{12}x^2\over \cancel3x^5}\]

OpenStudy (paxpolaris):

\[={6\over x^\left( {9+5-6} \right)}={6\over x^8}\]